A staircase contains three steps is 10 cm and 20 cm wide what should be the minimum horizontal velocity of a ball rolling of the uppermost plane so as to be directly the lower brain

Grade:11

1 Answers

Arun
5 years ago
o hit directly the lowest plane the ball has to cross the point B.
From A to B, the  ball has to travel to horizontal distance equal 40cm =40m
and vertical distance of 20cm=0.2m.
Let the time taken to reach point be t.u=0 m/s
h=0.2m
from the equation, h=ut+1/2gt²
we get
0.2=0+1/2 x9.8 t²
t²=0.4/9.8
t²=0.04
t=0.2
Let v be the horizontal velocity of ball than the horizontal distance travelled by the ball=0.2v
in order to clear the point B it must be equal to 0.40.
2v=0.4
v=2m/s
∴So minimum velocity required to to hit the lowest floor  is 2 m/s.